OpenStudy (anonymous):

METALS WILL BE GIVEN. Solve. Show all your work and explain your steps. the triagle on the left has a perimeter of 14. the triangle on the right has a perimeter of 21. what are x & y ?

OpenStudy (anonymous):



OpenStudy (anonymous):

Hmmm. So perimeter is all sides added together, can you write an equation for the perimeter of the first one?

OpenStudy (anonymous):

p = 2y + 1x is that right?

OpenStudy (anonymous):

Yep but you know the perimeter so go ahead and put that in to x+2y=14 now the second one

OpenStudy (anonymous):

3x + 5/4y = 21 is that right?

OpenStudy (anonymous):

Forgot a side 3x+(5/4)y+(5/4)y=21 which simplifies to 3x+(5/2)y=21 so now we have two equations 3x+(5/2)y=21 x+2y=14 any idea how to solve?

OpenStudy (anonymous):

We could use the elimination process right?

OpenStudy (anonymous):

You could. Multiply the second equation by -3 and then add the two equations together